### v2

# Financial sentiment analysis using FinBERT with application in predicting stock movement

Abstract
In this study, we integrate sentiment analysis within a financial framework by leveraging FinBERT, a fine-tuned BERT model specialized for financial text, to construct an advanced deep learning model based on Long Short-Term Memory (LSTM) networks. Our objective is to forecast financial market trends with greater accuracy. To evaluate our model’s predictive capabilities, we apply it to a comprehensive dataset of stock market news and perform a comparative analysis against standard BERT, standalone LSTM, and the traditional ARIMA models. Our findings indicate that incorporating sentiment analysis significantly enhances the model’s ability to anticipate market fluctuations. Furthermore, we propose a suite of optimization techniques aimed at refining the model’s performance, paving the way for more robust and reliable market prediction tools in the field of AI-driven finance.
1 Introduction
BERT (Bi-directional Encoder Representation for Transformers)Devlin et al. (2019) is a contextual representation language model been proved to be a very successful various natural language tasks. There has been various variants of BERT which focu on different tasks in NLP and different domain of language models. FinBERT, which is introduced in Araci (2019), is the first contextual pretrained language models which is trained specifically on a large scale of financial communication corpora. This is originally designed for the purpose of financial sentiment analysis especially focus on the need of capital market practitioners and researchers. There has been various applications of FinBERT in research, for example, Liu et al. (2020) studies financial text mining, Y Yang et al. (2020) studies financial communications, Huang et al. (2022) studies information extraction in finance etc.
Our main task in this paper is to use sentiment-based information in predicting the movement of the financial market. Financial data is usually presented in time series data, for example, stock prices. Traditional statistical method such as ARIMA has been used widely in the modelling of financial markets. However, statistical models like ARIMA does not capture well the nonlinearity in the financial data. People moved to deep learning methods especially in LSTM in studying financial time series, for example in Fjellstrom (2022). Our basic hypothesis is that the financial sentiment by news, social media, etc. have strong correlations to the movement of financial markets (stock market, foreign exchange, bond, etc.). This hypothesis is based on Efficient Market Hypothesis (EMH), which states that share prices reflect all information. Hence we are able to predict financial market using all those information. In particular, we believe that news contain a large amount of information. This hypothesis is a well-known, but the how to build quantitative models and making the prediction is still a challenging problem. We propose a model which takes advantage of Transformer based contextual language model for analyzing the sentiment in financial markets, and then using LSTM-based deep learning model in the prediction of the exact movement of the financial market by the sentiment. One thing need to be pointed our that, we are not doing casual analysis here, as market could influence news and news might also influence the market. We just study the relation of these two objects, and quantify this correlation.
We find the the LSTM model in general perform much better than ARIMA model, and sentiment analysis using FinBERT out perform LSTM, whereras FinBERT combined with LSTM acheived the best performance. This result complies with our assumption and anticipation.
2 Related Work
Transformer-based attention network for stock move- ment predictionZhang et al. (2022) proposed a novel model Transformer Encoder-based Attention Network (TEANet) framework for the prediction of stock price movement, and apply it to a trading simulation to test the validity of this model. Their model does not use sentiment analysis, and uses the transformer model for the time series data modelling, which is different than our sentiment-LSTM structure. However, their model would be helpful in our future work that we can replace the LSTM model by the TEANet for the sequencing modelling.
Bert-based financial sen- timent index and lstm-based stock return predictabilityHiew et al. (2019) studied similar problem as us, but with the base BERT for their sentiment analysis part. They proposed Financial Sentiment Index, which incorporated both the contextual channel and two other type of information sources which are option-implied sentiment based on the risk-neutral implied skewness, and the market implied sentiment based on market data. These three channel are then put together for the prediction of stock return. Our sentiment analysis is just using the raw sentiment information from the FinBERT, which might be tuned to incorporate Financial Sentiment Index in the future work. Their prediction model is also based on LSTM, which is then compared to a more traditional econometric model called VAR (Vector Autoregression). The VAR required lots of assumptions, while the noval model does not require. For the results, LSTM performed much better than the VAR which implied that LSTM captured nonlinearlity of the market information. Our model uses FinBERT in the sentiment analysis part, which is supposed to have much better performance in financial data sets. Hence we anticipate our model would our perform using base BERT.
Traditionally, the ARIMA model is used in modelling financial time series. This method does not capture well the nonlinearity of the information. Attention-based cnn-lstm and xgboost hybrid model for stock predictionShi et al. (2022) proposed a mixed model which combines the attention-based-covolutional neural network, LTSM, and then fine-tune by XGBoost. This paper did not use sentiment analysis. First, the model uses the traditional ARIMA for preprocessing the data. Then it uses attention-based CNN to extract deep features of the original stock data. After that, LSTM is used to mine the long term time series features. Finally, XGBoost is used to fine-tune the model performance. It turned out that attention-based CNN with Bi-LSTM achieved the the best performance across all base line models as well as single CNN or LSTM or BiLSTM. The model is this paper is similar to the sequencing model part of ours but with more stacking of various different models. Our model will not use the fusion model in the times series part, but in the future work we might adapt their ideas and method in improving the performance.
3 Data
We use the following data sets and data resources:
  • Pre-trained models, We use base BERT uncased and FinBERT from the Huggingface service.
  • Historical financial news archive, which contains news archive of more than 800 US companies for 12 years (up to 2019). This dataset contains date, ticker (of the stock), title, body, which are exactly information that we want. There has been people using this data set doing sentiment analysis. Hence we believe this would be a good start for our research and we decide to use it as our main dataset. By joining with the market data time series, we are able to perform sentiment analysis as well as the market prediction. This dataset is publicly available on Kaggle.com.
  • Daily Financial News for 6000+ Stocks. Similar to previous one, with more data but less information for each data (for example, only include header). This will complements our main dataset. This dataset is also available on Kaggle.com.
  • We also get stock market data from the yfinance API. For each ticker in previous datasets, we can get stock market information such as open price, closed price, percentage change, etc. This would be one of the main data besided sentiment in our LSTM model.
We use Historical financial news archive as our main dataset. We also merge the analyst_ratings_processed.csv from Daily Financial News for 6000+ Stocks and drop irrelevant columns. We have 1,056,471 rows of data in total, which consists of title, date, and ticker information of stocks.
Figure 1: A plot of MSFT closing price
4 The Model
Models Training Loss Val loss
ARIMA N/A 0.00975
LSTM 1.0206e-04 1.1937e-04
BERT + LSTM 1.0216e-04 3.3132e-04
FinBERT + DNN 1.4674e-04 3.4177e-04
FinBERT + LSTM 9.1701e-05 3.1975e-04
Table 1: Training and Testing loss on different models
Our model consists of two parts:
  1. 1.
    Financial Sentiment Analysis: We use the pre-trained FinBERT model in sentiment analysis, which is trained on a subset of Reuters TRC2 dataset and Financial PhraseBank from Malo et al. (2014). We preprocessing the We get the sentiment score from each financial news data for specific company by tickers. The result of the sentiment analysi would be one of the main data for the prediction of market movement.
  2. 2.
    LSTM for modelling financial times series data: after the sentiment analysis, we join the data of stock market information, which mainly consists of price information. Then we build sequence model using LSTM. Our LSTM module consists of the following layers: 1) a LSTM layer with 100 network units. We set the return_sequence to true so that the output of the layer will be another sequence of the same length. 2) Then we add another LSTM layer with 100 network units. Here we set the return_sequence to false for this time to only return the last output in the output sequence. 3) dense layer with 25 units. 4) Finally a dense layer with 1-dimensional output. We set the monitor the performance of the training by looking at validation loss, where we set 10% as validation set. The reason we choose small validation is that for financial data, the uncertainty plays an important role, we cannot forecast the future accurately especially in the long run.
We compare our model performance to the following baseline models:
  1. 1.
    Simple ARIMA model.
  2. 2.
    LSTM on both textual and numerical data.
  3. 3.
    FinBERT with DNN.
  4. 4.
    BERT with LSTM.
Figure 2: A plot of rolling mean and standard deviation for MSFT closing price
We use Min Square Error (MSE) as our main metric in both training and validation process since our target variable (close price) is numerical, and our model is a time series model. For ARIMA, we also use Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) during the tuning process.
5 Experiments
We started from the sentiment analysis using the FinBERT. Using the HuggingFace API, we are able to get sentiment from each financial news. We proposed a numerical sentiment index based on the price movement, which is defined as $NSI=1$ if $return>s$ and $NSI=0$ if $-s<return<s$ and $NSI=-1$ if $return<-s$. Here return is defined as
$return=\frac{ClosePrice_{t}-OpenPrice_{t}}{OpenPrice_{t}}$ (1)
the $t$ is the date and $s=0.01$ is our chosen threshold. Note that we can easily generalize this to the $k$-day return as
$return=\frac{ClosePrice_{t+k}-OpenPrice_{t}}{OpenPrice_{t}}$ (2)
NSI will be useful in fine-tune pre-trained FinBERT and study the correlation of textual sentiment and the price-based sentiment. Our task is to predict close price. In order to avoid the long-run inflation and size of different stocks, we use MinMax scalar before the model, and normalize the training variables. We trained our model with 90-10 split based on date, i.e. the first 90% days as training data and the remaining as testing data. The reason is the it is very hard to predict long term behavior of the Financial market, as there are too many factors can infuence the fluctuations of Financial markets, and some of them might be completely unexpected, for example, COVID-19, war, etc. We actually started from 80-20 splits, and the result was not so satisfying, which did not outperform a lot than our baseline LSTM model.
6 Analysis
From the result we see that our FinBERT sentiment analysis + LSTM model acheives the best training and validation loss. Base BERT + LSTM follows closely with it, which implies that our model has not fully taken advantage of the FinBERT language model since we only use the raw sentiment scores. In this paper, we only look at the sentiment from the titles of news, which contain of relatively small amount information and sentiment. This suggest us also trying to see other informations, for example, social network data like twitter, weibo, etc.
We also see that ARIMA actually capture the long run trend well, whereas our model is able to fit the short run trend which corresponding to the movement in short amount of the time pretty good. For a very short training time compared to other deep learning based models, AIRMA achieves MSE: 0.00975, MAE: 0.07154 and RMSE: 0.09876. As we see from the plot of predicted close price our model tend to under predict the close price in the long run. This suggest that in the future work, we might adopt similar method as in Shi et al. (2022), which preprocesses the data by ARIMA first, and then feed it into the deep neural network.
During the training process, we also see the the loss dropped significantly even just over 3 epochs, which is due to small size of the data. Note that even the whole training set is small, we only trained and tested by individual stock, and each stock only contains several thousands data on average. For example, the raw data for MSFT consists of 7875 rows. After the data processing, we have even less with 1440 rows left. This is challeging for deep learning based method since the model tends to overfit, which means that the model might capture more about the noise in the market, rather than the effective information. As we know that in stock market and even in financial markets in general, there has been strong correlations everywhere. We need to find a way to figure out these correlation, so that we might be able to put all stock data together and make a combined model.
Figure 3: Plot of predicted close price and real close price
7 Conclusion
In this paper, we presented a model that leverages financial sentiment analysis to enhance the prediction of stock market movements. Our approach combined the specialized FinBERT, a transformer-based language model trained on financial corpora, with LSTM networks to interpret and utilize time-series data effectively. We demonstrated that this fusion outperforms traditional models such as ARIMA, as well as LSTM models without sentiment analysis, in capturing the nuances of market fluctuations.
Despite the encouraging results, our study identified several areas for improvement and further research:
  1. 1.
    The model exhibits a tendency to underpredict long-term trends. This suggests a potential integration of ARIMA for preprocessing to improve long-range forecasts, as demonstrated by other hybrid approaches in the literature.
  2. 2.
    Limitations in the size of the dataset for individual stocks led to concerns about overfitting. Future work could explore the inclusion of a broader array of financial news or social media data to enrich the training set and mitigate this issue.
  3. 3.
    While FinBERT’s sentiment analysis showed advantages over base BERT, the margin of improvement was modest. Given the findings in ”Is Domain Adaptation Worth Your Investment? Comparing BERT and FinBERT on Financial Tasks” Peng et al. (2021), it would be prudent to consider further fine-tuning of FinBERT on a larger or augmented dataset, potentially amplifying the benefits of domain-specific modeling.
In conclusion, our sentiment-based stock market prediction model represents a promising step forward in the application of AI to financial forecasting. The interplay between news sentiment and market behavior, while complex, can be harnessed to produce more accurate and reliable predictions. As we continue to refine our model and incorporate a wider spectrum of data sources, we anticipate further improvements in performance and a deeper understanding of the intricate dynamics at play in financial markets.